OpenStudy (anonymous):

Solve the following system of equations: -2x + y = 1 -4x + y = -1 A. (3, 1) B(-1, 3) C. (-1, -3) D. (1, 3)

OpenStudy (anonymous):

D

OpenStudy (anonymous):

All you would do is plug in the coordinates of each answer to both of the problems.

OpenStudy (radar):

Isolate y in each equation getting: y=2x + 1 and y=4x -1 We know y=y so then the 2x +1 = 4x -1 solve for x getting 2x = 2 or x =1. Now substitute 1 for x in either of the original equation and solve for y.